OpenStudy (anonymous):

if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8

OpenStudy (anonymous):

−8 + y −8 − y 8 + y 8 − y

jimthompson5910 (jim_thompson5910):

start with the second equation −x − y = −8 and add y to both sides to get −x − y = −8 −x − y+y = −8 +y -x + 0y = -8 + y -x = -8 + y Then multiply both sides by -1 to fully isolate x -x = -8 + y -1*(-x) = -1*(-8 + y) x = 8 - y

jimthompson5910 (jim_thompson5910):

alternatively, you can do it this way −x − y = −8 − y = −8 + x -y + 8 = x x = -y + 8 x = 8 - y